A 4500kg open train car is rolling on frictionless rails at 21.0m/s when it starts pouring rain. A few minutes later, the car's speed is 20.0 m/s. What mass of water has collected in the car?
Answer #1

Momentum before = MOmentum after
mv = mv

4500 x 22 = Y x 20
94500 = 20Y
Y = 4725kg

Mass of Rain = Mass of Train after - Mass of Train before = 4725-4500 = 225kg
Answer #2

4500*21 = (4500+m)*20

m = 225 Kg